# Sampling 3D Molecular Conformers with Diffusion Transformers

### v1

@@ Abstract
Diffusion Transformers (DiTs) have demonstrated strong performance in generative modeling, particularly in image synthesis, making them a compelling choice for molecular conformer generation. However, applying DiTs to molecules introduces novel challenges, such as integrating discrete molecular graph information with continuous 3D geometry, handling Euclidean symmetries, and designing conditioning mechanisms that generalize across molecules of varying sizes and structures. We propose DiTMC, a framework that adapts DiTs to address these challenges through a modular architecture that separates the processing of 3D coordinates from conditioning on atomic connectivity. To this end, we introduce two complementary graph-based conditioning strategies that integrate seamlessly with the DiT architecture. These are combined with different attention mechanisms, including both standard non-equivariant and SO(3)-equivariant formulations, enabling flexible control over the trade-off between between accuracy and computational efficiency. Experiments on standard conformer generation benchmarks (GEOM-QM9, -DRUGS, -XL) demonstrate that DiTMC achieves state-of-the-art precision and physical validity. Our results highlight how architectural choices and symmetry priors affect sample quality and efficiency, suggesting promising directions for large-scale generative modeling of molecular structures. Code available at https://github.com/ML4MolSim/dit_mc.
Keywords Generative Modeling  $\cdot$ Flow Matching  $\cdot$ Diffusion Transformers  $\cdot$ Geometric Deep Learning  $\cdot$ Molecular Simulation  $\cdot$ Conformer Generation
@@ 1 Introduction
Recent years have seen significant progress, enabled by the development of specialized architectures for the generation of molecules [1, 2, 3, 4, 5, 6, 7, 8, 9] and materials [10, 11, 12]. This is in contrast to image and video synthesis, where the more generalized diffusion transformer (DiT) architecture [13] has become a leading model, consistently delivering strong performance and efficiency across diverse applications [14, 15, 16, 17]. Adapting DiTs, which were originally developed for grid-structured image data, to continuous, irregular molecular geometries poses unique challenges, which need to be addressed to unlock the potential of this powerful architecture for molecular conformer generation and leverage prior advances from vision research. Key design questions include how to encode molecular connectivity and incorporate Euclidean symmetries, such as translational and rotational invariance/equivariance.
In this work, we address these conceptual challenges and propose DiTMC, a new DiT-style architecture for molecular conformer generation. We introduce novel conditioning strategies based on molecular graphs, enabling 3D conformer generation. Our modular architectural design allows us to systematically investigate the impact of different self-attention mechanisms within the DiT architecture on conformer generation quality and efficiency. We conduct a comparative study including standard (non-equivariant) self-attention with both absolute and relative positional embeddings and an explicitly $\mathrm{SO}(3)$-equivariant variant. While exact equivariance can positively impact performance, it also incurs significant computational costs. We find that simpler attention mechanisms are highly scalable and still perform competitively. Multiple of the tested DiT variants achieve state-of-the-art (SOTA) precision on established conformer generation benchmarks. Further, the ensembles of molecular structures generated by our models exhibit increased realism, as evidenced by the high accuracy of physical properties extracted from them. To summarize, our work contains the following main contributions:
  • We propose two complementary conditioning strategies based on trainable conditioning tokens for (pairs of) atoms extracted from molecular graphs, which are designed to align with the architectural principles of DiTs.
@@
  • Based on our insights, we present a simple, non-equivariant, yet expressive DiT architecture that achieves state-of-the-art (SOTA) precision and physical validity on established benchmarks. Its performance improves with model scaling, making it a promising candidate for large-scale molecular conformer generation.
@@ 2 Related Work
Generative Modeling
Generative models create diverse, high-quality samples from an unknown data distribution. They are widely used in image generation [18, 19], text synthesis [20] and the natural sciences [21, 22, 23]. Most recent approaches learn a probabilistic path from a simple prior to the data distribution, enabling both efficient sampling and likelihood estimation [24, 25, 26, 27, 28]. This is achieved by different modeling paradigms, including flow matching [26] or denoising diffusion [24].
(Diffusion) Transformers
Originally introduced for natural language processing [29], transformer architectures have become SOTA in domains such as computer vision [18], and recently also found widespread adoption in quantum chemistry, e.g., for protein prediction [30, 31], 3D molecular generation [32, 33] or molecular dynamics simulations [34, 35, 36]. In the context of generative modeling, diffusion transformers [13] (DiTs) have emerged as powerful tools incorporating conditioning tokens, e.g., to prompt image generation [13] or to design molecules and materials with desirable properties [37]. This work applies prototypical DiTs to molecular conformer generation.
Molecular Conformer Generation
Molecular conformer generation aims to find atomic arrangements (Cartesian coordinates) consistent with a given molecular graph. Numerous ML approaches have been proposed for sampling conformers, all aiming to improve upon conventional methods. While early approaches were based on RDKit [38] or variational auto encoders [39, 40, 41], more recent advances employ diffusion and flow-based models [42, 43], including E-NFs [9], CGCF [44], GeoDiff [45], TorsionDiff [46], MCF [47], ET-Flow [32], and DMT [48].
@@ 3.1 Molecular Conformers
A molecule can be associated with a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where the nodes $\mathcal{V}$ correspond to atoms and the edges $\mathcal{E}$ represent chemical bonds between them. The nodes and edges contain information about their types, bond orders, and additional structural features such as branches, rings, and stereochemistry. What is missing is the exact spatial arrangement of the $N=\lvert\mathcal{V}\rvert$ atoms as a 3D point cloud $\boldsymbol{x}\in\mathbb{R}^{N\times 3}$ in Euclidean space. Only the relative distances between atoms are relevant, as translating or rotating the entire point cloud $\boldsymbol{x}$ does not change the identity of the conformer. We frame conformer prediction as sampling from the conditional probability distribution $p(\boldsymbol{x}\,|\,\mathcal{G})$, which will guide the design of our model architectures (Sec. 4).
@@ 3.2 Conditional Flow Matching
Starting from an easy-to-sample base distribution $q_{0}:\mathbb{R}^{d}\mapsto\mathbb{R}_{\geq 0}$, a generative process creates samples from a target distribution $q_{1}:\mathbb{R}^{d}\mapsto\mathbb{R}_{\geq 0}$ [27]. Here, $q_{1}$ models the molecular conformer data with $d=N\times 3$. We aim to learn a time-dependent vector field $u_{t}(\boldsymbol{x}):[0,1]\times\mathbb{R}^{N\times 3}\mapsto\mathbb{R}^{N\times 3}$, which defines an ordinary differential equation (ODE) whose solution pushes samples $\boldsymbol{x}_{0}\in\mathbb{R}^{N\times 3}$ from the prior to samples $\boldsymbol{x}_{1}\in\mathbb{R}^{N\times 3}$ from the data distribution.
We describe this transformation in terms of a stochastic interpolant $\boldsymbol{x}_{t}$ [49, 26, 50]. A noisy sample at time $t\in[0,1]$ is defined as
@@
where $\boldsymbol{\epsilon}\in\mathbb{R}^{N\times 3}$ is drawn from the standard normal distribution $\mathcal{N}(0,\boldsymbol{I})$ and scaled by a constant $\sigma\in\mathbb{R}_{\geq 0}$. We remark that $t$ represents progress along this interpolation path, not physical time. Notably, stochastic interpolants enable transformations between arbitrary distributions and allow us to assess the performance of the generative process under varying prior distributions $q_{0}$. This contrasts with, e.g., score based diffusion methods [24, 51], which typically assume an isotropic Gaussian prior.
@@
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}(0,1),\boldsymbol{x}\sim p_{t}(\boldsymbol{x})}\ \Big{|}\Big{|}u_{t}(\boldsymbol{x})-\boldsymbol{v}^{\theta}(\boldsymbol{x},t)\Big{|}\Big{|}.$ (3)
@@
However, for arbitrary distributions $q_{0}$ and $q_{1}$, the objective in Eq. 3 is computationally intractable [52]. Instead, we consider the expectation over interpolated point pairs from the two distributions. Eq. 1 defines a conditional probability distribution $p_{t}(\boldsymbol{x}|\boldsymbol{x}_{0},\boldsymbol{x}_{1})=\mathcal{N}(\boldsymbol{x}|(1-t)\cdot\boldsymbol{x}_{0}+t\cdot\boldsymbol{x}_{1},\sigma^{2})$, with conditional vector field $\boldsymbol{u}_{t}(\boldsymbol{x}|\boldsymbol{x}_{0},\boldsymbol{x}_{1})=\boldsymbol{x}_{1}-\boldsymbol{x}_{0}$ [27]. The ability to directly sample from the conditional probability via Eq. 1 allows formulating the conditional flow matching (CFM) objective
@@
As shown in Ref. [26], the gradients of the two losses coincide, $\nabla_{\theta}\mathcal{L}_{\text{FM}}=\nabla_{\theta}\mathcal{L}_{\text{CFM}}$, thereby recovering the vector field that defines the probability flow ODE in Eq. 2. Following prior work [53, 37], we reparametrize the training objective to predict noise-free data $\boldsymbol{x}_{0}^{\theta}$ directly. During inference, we invert the reparametrization to obtain $\boldsymbol{v}^{\theta}$ for sampling (see Appendix C).
@@ 4 A New Diffusion Transformer for Molecular Conformer Sampling
We now describe the methodological advances of our work. We propose DiTMC, a new DiT-style architecture for conformer generation by learning a vector field using the loss in Eq. 4. As outlined in Sec. 3.1, this involves sampling from a conditional probability $p(\boldsymbol{x}|\mathcal{G})$, where $\mathcal{G}$ is the molecular graph representing atomic connectivity. Therefore, we choose our model to be a function $\boldsymbol{v}^{\theta}(\boldsymbol{x},t,\mathcal{G})$.
We begin by initializing the node features using learnable embeddings $\boldsymbol{e}(z_{i})\in\mathbb{R}^{H}$, where $z_{i}\in\mathbb{N}_{+}$ denotes the atomic number of atom $i$ [35]. These embeddings are then processed through multiple DiT blocks. Each DiT block transforms the current node features (input tokens) =$\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\boldsymbol{h}_{i}\in\mathbb{R}^{H}\}$ to a set of output tokens $\mathcal{H}^{\prime}$, which serve as input for the next block. Additional inputs consist of positional embeddings $\mathcal{P}$ (representing the position of nodes in Euclidean space), as well as conditioning tokens $\mathcal{C}$, for example for encoding the time ($\boldsymbol{c}^{t}$) and graph structure ($\boldsymbol{c}^{\mathcal{G}}$).
The central operation in each DiT block is a self-attention update,
$\boldsymbol{h}_{i}=\boldsymbol{h}_{i}+\text{ATT}(\mathcal{H},\mathcal{P},\mathcal{C}^{\text{Pair}})_{i}\,,$ (5)
followed by a node-wise Multi-Layer Perceptron (MLP). Here, $\mathcal{C}^{\text{Pair}}$ are special conditioning tokens acting on pairs of atoms (see Sec. 4.1). Similar to applications of DiTs in image synthesis, adaptive layer norm (AdaLN) and adaptive scale (AdaScale) [13] are used for conditioning (see Fig. 1A) based on per-atom bias and scaling parameters
$\alpha_{1i},\beta_{1i},\gamma_{1i},\alpha_{2i},\beta_{2i},\gamma_{2i}=\text{MLP}(\boldsymbol{c}^{t}+\boldsymbol{c}_{i}^{\mathcal{G}})\,.$ (6)
The final output is a 3D vector (“velocity”) per atom, which is extracted from a readout layer. The whole architecture is summarized in Fig. 1A. In the following, we describe the individual components of the architecture in more detail. Training details are given in Appendix A.
4.1 Conditioning
Each DiT block receives conditioning signals that encode information about the time step $t$, as well as atom- and pair-wise information derived from a molecular graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. Pair-wise information is used during the self-attention update (see Eq. 5 and Sec. 4.3), whereas atom-wise information and the time conditioning signal are injected via adaptive layer norms (see Eq. 6).
Time conditioning
The current time $t$ of the latent state $\boldsymbol{x}_{t}$ is encoded via a two-layer MLP as
@@ 4.1 Conditioning
Atom-wise conditioning
Atom-wise conditioning tokens are obtained from a GNN inspired by the processor module of the MeshGraphNet (MGN) framework [54]
@@
where $\boldsymbol{c}_{i}^{\mathcal{G}}$ denotes the final node representation for atom $i$. See Appendix I for details on the GNN.
Pair-wise conditioning
In addition to node representations, the GNN also produces edge-level representations, which can be used to define pair-wise conditioning tokens:
[…]
These tokens only capture interactions between bonded atoms, i.e., when $(i,j)\in\mathcal{E}$. Conditioning tokens for non-bonded pairs are set to a learnable vector $\boldsymbol{\bar{c}}^{\mathcal{G}}$. Since self-attention operates on all atom pairs $(i,j)$, even if they are not connected by a chemical bond, we also define an alternative pair-wise conditioning method inspired by the Graphormer architecture [55]
[…]
where $s(i,j)$ denotes the graph geodesic (i.e., the shortest path between atoms $i$ and $j$ via edges in $\mathcal{G}$). This formulation allows conditioning on all atom pairs, even if they are not directly connected. In the following, we refer to these two variants as bond-pair (Eq. 9) and all-pair (Eq. 10) conditioning, respectively (see Sec. 5 for empirical comparisons between both methods).
4.2 Positional Embeddings ( PE )
We inject information about the current atomic positions =$\mathcal{R}=\{\vec{r}_{1},\dots\vec{r}_{N}\,|\,\vec{r}_{i}\in\mathbb{R}^{3}\}$ of the latent state $\boldsymbol{x}_{t}$ via positional embeddings (PEs). To that end, we assume the positions to have zero center of mass (see section A.1). As mentioned in Sec. 3.1, conformer identity is determined only by the relative positions of atoms (i.e., it is invariant under rigid translations and rotations), but ensuring that these symmetries are preserved within the network often incurs computational overhead.
Therefore, we examine a representative range of PEs that vary in the number of Euclidean symmetries they respect by construction, which affects how the latent representations transform under translations and rotations. This includes absolute and relative PEs, motivated by a similar distinction made in other application domains [56, 57]. Inspired by recent advances in geometric deep learning [58, 59, 60], we also propose a PE strategy, obeying all Euclidean symmetries. Our aim is to find an acceptable tradeoff between model accuracy, efficiency and scalability, by scanning the design space of PE strategies.
Absolute Positional Embeddings ( aPE )
Following Refs. [31, 37] they are calculated as

### v2

@@ Abstract
Diffusion Transformers (DiTs) have demonstrated strong performance in generative modeling, particularly in image synthesis, making them a compelling choice for molecular conformer generation. However, applying DiTs to molecules introduces novel challenges, such as integrating discrete molecular graph information with continuous 3D geometry, handling Euclidean symmetries, and designing conditioning mechanisms that generalize across molecules of varying sizes and structures. We propose DiTMC, a framework that adapts DiTs to address these challenges through a modular architecture that separates the processing of 3D coordinates from conditioning on atomic connectivity. To this end, we introduce two complementary graph-based conditioning strategies that integrate seamlessly with the DiT architecture. These are combined with different attention mechanisms, including both standard non-equivariant and SO(3)-equivariant formulations, enabling flexible control over the trade-off between between accuracy and computational efficiency. Experiments on standard conformer generation benchmarks (GEOM-QM9, -DRUGS, -XL) demonstrate that DiTMC achieves state-of-the-art precision and physical validity. Our results highlight how architectural choices and symmetry priors affect sample quality and efficiency, suggesting promising directions for large-scale generative modeling of molecular structures. Code is available at https://github.com/ML4MolSim/dit_mc.
@@ 1 Introduction
Recent years have seen significant progress, enabled by the development of specialized architectures for the generation of molecules [gebauer2019symmetry, masuda2020generating, ragoza2020learning, gebauer2022inverse, luo20213d, simonovsky2018graphvae, de2018molgan, hoogeboom2022equivariant, garcia2021n] and materials [jiao2023crystal, vecchio2024matfuse, zeni2025generative]. This is in contrast to image and video synthesis, where the more generalized diffusion transformer (DiT) architecture [peebles2023DiT] has become a leading model, consistently delivering strong performance and efficiency across diverse applications [bao2023all, chen2024gentron, liu2024sora, ma2025latte]. Adapting DiTs, which were originally developed for grid-structured image data, to continuous, irregular molecular geometries poses unique challenges, which need to be addressed to unlock the potential of this powerful architecture for molecular conformer generation. Key design questions include how to encode molecular connectivity and incorporate Euclidean symmetries, such as translational and rotational invariance/equivariance.
In this work, we address these conceptual challenges and propose DiTMC, a new DiT-style architecture for molecular conformer generation. We introduce novel conditioning strategies based on molecular graphs, enabling the generation of 3D structures. Our modular architectural design allows us to systematically investigate the impact of different self-attention mechanisms within the DiT architecture on conformer generation quality and efficiency. We conduct a comparative study including standard (non-equivariant) self-attention with both absolute and relative positional embeddings and an explicitly $\mathrm{SO}(3)$-equivariant variant. While exact equivariance can positively impact performance, it also incurs significant computational costs. We find that simpler attention mechanisms are highly scalable and still perform competitively. Multiple of the tested DiTMC variants achieve state-of-the-art precision on established conformer generation benchmarks. Moreover, the molecular structure ensembles generated by our models align more closely with physical reality, as evidenced by the high accuracy of physical properties extracted from them. To summarize, our work contains the following main contributions:
  • We propose two complementary conditioning strategies based on trainable conditioning tokens for (pairs of) atoms extracted from molecular graphs, which are designed to align with the architectural principles of DiTs. We propose to condition our self-attention formulation on geodesic graph distances extracted from molecular graphs and demonstrate that it significantly increases performance of our model.
@@
  • Based on our insights, we present a simple, non-equivariant, yet expressive DiT architecture that achieves state-of-the-art precision and physical validity on established benchmarks. Its performance improves with model scaling, making it a promising candidate for large-scale molecular conformer generation.
@@ 2 Related Work
Generative Modeling Generative models create diverse, high-quality samples from an unknown data distribution. They are widely used in image generation [dosovitskiy2021an, rombach2022high], text synthesis [gu2022vector] and the natural sciences [jumper2021highly, alakhdar2024diffusion, hsu2024score]. Most recent approaches learn a probabilistic path from a simple prior to the data distribution, enabling both efficient sampling and likelihood estimation [ho2020denoising, kingma2021variational, lipman2023flow, albergo2023stochastic, kingma2023understanding]. This is achieved by different modeling paradigms, including flow matching  [lipman2023flow] or denoising diffusion [ho2020denoising].
(Diffusion) Transformers Originally introduced for natural language processing [vaswani2017attention], transformer architectures have become state-of-the-art in computer vision [dosovitskiy2021an], and recently also found widespread adoption in quantum chemistry, e.g., for protein prediction [abramson2024accurate, geffner2025proteina], 3D molecular generation [hassan2024etflow, lewis2024scalable] or molecular dynamics simulations [tholke2022equivariant, frank2024fastandstable, liao2024equiformerv]. In the context of generative modeling, diffusion transformers [peebles2023DiT] (DiTs) have emerged as powerful tools incorporating conditioning tokens, e.g., to prompt image generation [peebles2023DiT] or to design molecules and materials with desirable properties [joshi2025all]. This work applies prototypical DiTs to molecular conformer generation.
Molecular Conformer Generation Molecular conformer generation aims to find atomic arrangements (Cartesian coordinates) consistent with a given molecular graph. Numerous ML approaches have been proposed for sampling conformers, all aiming to improve upon conventional methods.
While early approaches were based on RDKit [riniker2015better] or variational auto encoders [mansimov2019molecular, simm2019generative, xu2021end], more recent advances employ diffusion and flow-based models [luo2021predicting, shi2021learning], including E-NFs [garcia2021n], CGCF [xu2021learning], GeoDiff [xu2022geodiff], TorsionDiff [jing2022torsional], MCF [wang2023swallowing], ET-Flow [hassan2024etflow], and DMT [liu2025nextmol].
De-Novo Molecular Generation Instead of conditioning on a given molecular graph, several existing approaches tackle the problem of finding novel and stable molecules given only a set of atoms [huang2023mdm, hoogeboom2022equivariant, xu2023geometric]. Bridging the gap between unconditional and conformer generation, universal models have been proposed that can solve a multitude of tasks. In particular, models within the Uni-Mol model family [zhou2023uni, ji2024uni], make use of graph geodesic distances. However, their design requires tracking high-dimensional pair-representations throughout the architecture, which makes it less efficient.
@@ 3.1 Molecular Conformers
A molecule can be represented as a molecular graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the nodes $\mathcal{V}$ correspond to atoms and the edges $\mathcal{E}$ represent chemical bonds between them. The nodes and edges contain information about their types, bond orders, and additional structural features such as branches, rings, and stereochemistry. The missing component is the exact spatial arrangement of the $N=\lvert\mathcal{V}\rvert$ atoms, represented as a 3D point cloud $\boldsymbol{x}\in\mathbb{R}^{N\times 3}$ in Euclidean space. Only the relative distances between atoms are relevant, as translating or rotating the entire point cloud $\boldsymbol{x}$ does not change the identity of the conformer. We frame conformer prediction as sampling from the $\mathrm{SE}(3)$-invariant conditional probability distribution $p(\boldsymbol{x}\,|\,\mathcal{G})$, which will guide the design of our model architectures in Sec. 4.
@@ 3.2 Conditional Flow Matching
Starting from an easy-to-sample base distribution $q_{0}:\mathbb{R}^{d}\mapsto\mathbb{R}_{\geq 0}$, a generative process creates samples from a target distribution $q_{1}:\mathbb{R}^{d}\mapsto\mathbb{R}_{\geq 0}$ [albergo2023stochastic]. Here, $q_{1}$ models the molecular conformer data with $d=N\times 3$. We aim to learn a time-dependent vector field $u_{t}(\boldsymbol{x}):[0,1]\times\mathbb{R}^{N\times 3}\mapsto\mathbb{R}^{N\times 3}$, which defines an ordinary differential equation (ODE) whose solution pushes samples $\boldsymbol{x}_{0}\in\mathbb{R}^{N\times 3}$ from the prior to samples $\boldsymbol{x}_{1}\in\mathbb{R}^{N\times 3}$ from the data distribution. We describe this transformation in terms of a stochastic interpolant $\boldsymbol{x}_{t}$ [albergo2023building, lipman2023flow, liu2023flow]. A noisy sample at time $t\in[0,1]$ is defined as
@@
where $\boldsymbol{\epsilon}\in\mathbb{R}^{N\times 3}$ is drawn from the standard normal distribution $\mathcal{N}(0,\boldsymbol{I})$ and scaled by a constant $\sigma\in\mathbb{R}_{\geq 0}$. We remark that $t$ represents progress along this interpolation path, not physical time. Notably, stochastic interpolants enable transformations between arbitrary distributions and allow us to assess the performance of the generative process under varying prior distributions $q_{0}$. This contrasts with, e.g., score based diffusion methods [ho2020denoising, song2021scorebased], which typically assume an isotropic Gaussian prior.
@@
[…]
@@
However, for arbitrary distributions $q_{0}$ and $q_{1}$, the objective in Eq. 3 is computationally intractable [tong2024improving]. Instead, we consider the expectation over interpolated point pairs from the two distributions. Eq. 1 defines a conditional probability distribution $p_{t}(\boldsymbol{x}|\boldsymbol{x}_{0},\boldsymbol{x}_{1})=\mathcal{N}(\boldsymbol{x}|(1-t)\cdot\boldsymbol{x}_{0}+t\cdot\boldsymbol{x}_{1},\sigma^{2})$, with conditional vector field $\boldsymbol{u}_{t}(\boldsymbol{x}|\boldsymbol{x}_{0},\boldsymbol{x}_{1})=\boldsymbol{x}_{1}-\boldsymbol{x}_{0}$ [albergo2023stochastic]. The ability to directly sample from the conditional probability via Eq. 1 allows formulating the conditional flow matching (CFM) objective
@@
As shown in Ref. [lipman2023flow], the gradients of the two losses coincide, $\nabla_{\theta}\mathcal{L}_{\text{FM}}=\nabla_{\theta}\mathcal{L}_{\text{CFM}}$, thereby recovering the vector field that defines the probability flow ODE in Eq. 2. Following prior work [yim2023fast, joshi2025all], we reparametrize the training objective to predict noise-free data $\boldsymbol{x}_{0}^{\theta}$ directly. During inference, we invert the reparametrization to obtain $\boldsymbol{v}^{\theta}$ for sampling (see Appendix C for details).
@@ 4 A New Diffusion Transformer for Molecular Conformer Sampling
We now describe the methodological advances of our work. We propose DiTMC, a new DiT-style architecture for conformer generation by learning a vector field using the loss in Eq. 4. As outlined in Sec. 3.1, this involves sampling from a conditional probability $p(\boldsymbol{x}\,|\,\mathcal{G})$, where $\mathcal{G}$ is the molecular graph representing atomic connectivity. Therefore, we choose our model to be a function $\boldsymbol{v}^{\theta}(\boldsymbol{x},t,\mathcal{G})$, where the final output is a 3D velocity vector per atom, which is extracted from a readout layer.
Next, we outline the key components of the DiTMC architecture, with an overview shown in Fig. 1A. Training and architectural details are provided in Appendix B and Appendix D, respectively.
4.1 Conditioning Tokens
We begin by defining the conditioning tokens in DiTMC. Each DiTMC block receives a time conditioning token $\boldsymbol{c}^{t}\in\mathbb{R}^{H}$, as well as atom-wise conditioning tokens $\mathcal{C}_{\text{atom}}^{\mathcal{G}}=\{\boldsymbol{c}_{i}^{\mathcal{G}}\in\mathbb{R}^{H}\,|\,i\in[N]\}$, and pair-wise conditioning tokens $\mathcal{C}_{\text{pair}}^{\mathcal{G}}=\{\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{H}\,|\,i,j\in[N]\}$ derived from $\mathcal{G}=(\mathcal{V},\mathcal{E})$.
Time conditioning The current time $t$ of the latent state $\boldsymbol{x}_{t}$ is encoded via a two-layer MLP as
@@ 4.1 Conditioning Tokens
Atom-wise conditioning Atom-wise graph conditioning tokens are obtained from a GNN inspired by the processor module of the MeshGraphNet (MGN) framework [pfaff2020learning] as
@@
where $\boldsymbol{c}_{i}^{\mathcal{G}}$ denotes the final node representation for atom $i$. See Appendix D.3 for details on the GNN.
Pair-wise conditioning We define pair-wise graph conditioning tokens inspired by the Graphormer architecture [ying2021graphormer] as
$\boldsymbol{c}_{ij}^{\mathcal{G}}=\text{MLP}\left(s(i,j)\right),$ (7)
where $s(i,j)$ denotes the graph geodesic, i.e., the shortest path between atoms $i$ and $j$ in $\mathcal{G}$. This formulation allows conditioning on all atom pairs, including those not directly connected by a bond.
4.2 Positional Embeddings
To encode the atomic positions $\mathcal{R}=\{\vec{r}_{1},\dots,\vec{r}_{N}\,|\,\vec{r}_{i}\in\mathbb{R}^{3}\}$ of the latent state $\boldsymbol{x}_{t}$, we use positional embeddings (PEs). We examine a representative range of positional embeddings that vary in the number of Euclidean symmetries they respect by construction, which affects how the latent representations transform under translations and rotations. We denote the set of positional embeddings by $\mathcal{P}$, where $\mathcal{P}=\{\boldsymbol{p}_{i}\,|\,i\in[N]\}$ for the atom-wise (aPE) embeddings, and $\mathcal{P}=\{\boldsymbol{p}_{ij}\,|\,i,j\in[N],\,i\neq j\}$ for pair-wise (rPE or PE(3)) embeddings. Without loss of generality, we assume that the positions are centered such that the center of mass vanishes (see Appendix A).
Absolute Positional Embeddings Following Refs. [geffner2025proteina, joshi2025all], atom-wise absolute Positional Embeddings (aPE) are calculated as